OpenStudy (anonymous):

-18-6k=6(1+3k) i need help solving after you -18-6k=6+18k -6 -6 -12-6k=18k am i on the right track?

OpenStudy (anonymous):

-18-6k=6(1+3k) -18-6k=6+18k -18-6=18k+6k -24=24k k=24/24 k=1

OpenStudy (anonymous):

yes you are dear :)

OpenStudy (anonymous):

i appreciate the help. im a little confused on the 3rd line

OpenStudy (anonymous):

ouh,u should put all number that have 'k' together :)

OpenStudy (anonymous):

got it ?

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

k is -1

OpenStudy (anonymous):

omg.. -1 :3 hehe..sorry